OpenStudy (anonymous):

What is the discriminant of 9x2 + 2 = 10x? A:–356 B:–172 C:28 D:72

OpenStudy (astrophysics):

@ParthKohli

Parth (parthkohli):

@SarthKohli

OpenStudy (astrophysics):

Discriminant is \[b^2-4ac\] where your equation is of the form \[ax^2+bx+c=0\]

MsBrains (ms-brains):

What do you think the answer is? @seanzilla

OpenStudy (triciaal):

welcome to openstudy!! follow the directions given by @Astrophysics above

OpenStudy (triciaal):

put in the general format identify your values for a, b, c substitute and compute

OpenStudy (astrophysics):

Maybe lets put your equation in a better form so we can compare it to \[ax^2+bx+c=0\] \[9x^2+2=10x \implies 9x^2-10x+2=0\] try now

OpenStudy (anonymous):

Oh okay i understand the answer is 28! C!

OpenStudy (astrophysics):

Well done! :)

Parth (parthkohli):

Nice work. Can you teach these to me afterwards @Astrophysics

OpenStudy (anonymous):

Why don't I get a MEDAL? HM???????????????????????????????????

ILovePuppiesLol (ilovepuppieslol):

cause u didn't help xD @browniecat

OpenStudy (anonymous):

lol thanks for the 3 medals

OpenStudy (anonymous):

Keep em comin

OpenStudy (bhumi273):

D=\[b ^{2}-4ac\]

OpenStudy (bhumi273):

\[9\times2+2-10x=0\]

OpenStudy (bhumi273):

here, a=18 b=-10x c=2

OpenStudy (bhumi273):

\[D= (-10x)^{2}-4(18)(2)\]

OpenStudy (bhumi273):

\[100x ^{2}-144\]

OpenStudy (bhumi273):

\[100x ^{2}=144\]

OpenStudy (bhumi273):

\[x^{2}=\frac{ 144 }{ 100 }\]

OpenStudy (bhumi273):

\[x=\sqrt{1.44}\]